a polynomial p(x) is such that p(0)=5 ,p(1)=4 ,p(2)=9,p(3)=20 ,the minimum degree it can have

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Plotting the four points, a parabola mightpass through them.
The general form of parabola is p(x)=ax2+bx+c
p(0)=c
P(1)=a+b+c=4
P(2)=4a+2b+c=9
P(3)=9a+3b+c=20
On solving the above equations we get
P(x)=3x2-4x+5=0
Hence the minimum degree it can have is 2
Thanks and Regards
Shaik Aasif
askIITians faculty